Management has been having a quality problem with one of its computer lines. Three samples of 40 computers each have been taken and recorded below. You have been asked to examine the following samples and their defects to determine the average percentage defective. Using a 95 percent confidence level, graph the control chart. Make some conclusions about the process.

SAMPLE             

                                            1              2              3             

Number of Errors              5              5              9                             
Answer #1

Answer: 95% confidence corresponds to z value of 1.96

Day Defectives Observations
1 5 40
2 5 40
3 9 40
total 19 120
steps and formulas, for z=1.96
Proportion of defects=Pbar total defectives/total observations 0.158
Q= 1-P 0.842
N= average sample size 40
Standard deviation squareroot(P*Q/N) 0.058
UCL= P + z*squareroot(P*Q/N) 0.271
LCL= P - z*squareroot(P*Q/N) 0.045

P chart

The process is in control: As we can see that NO sample is beyond control limits in charts

use the table below

Day Defectives Observations proportion of defect=defectives/observations P UCL LCL
1 5 40 0.125 0.158 0.271 0.045
2 5 40 0.125 0.158 0.271 0.045
3 9 40 0.225 0.158 0.271 0.045